OpenStudy (sidsiddhartha):

INTEGRATE: \[\Large \int\limits_{-\infty}^{\infty}\frac{ 1 }{ x^4+1 }dx\]

OpenStudy (sidsiddhartha):

which technique i should use Residue? or something else

OpenStudy (anonymous):

Residue would be the simplest way, yes.

OpenStudy (sidsiddhartha):

yeah but it is consuming much time :(

OpenStudy (anonymous):

Factor the denominator (so you'll have to find the fourth roots of -1), then compute. The integral will be equivalent to \[2\pi i\left(\large\sum_{\text{Im}z>0}\text{Res}f(z)+\frac{1}{2}\sum_{\text{Im}z=0}\text{Res}f(z)\right)\]

OpenStudy (sidsiddhartha):

i'm considering \[\int\limits_{C}^{}f(z)dz\] so \[f(z)=\frac{ 1 }{ z^4+1 }\] which gives me \[\Large z^4=-1\rightarrow z=[\cos(2n+1)\pi+i \sin(2n+1)\pi]^{1/4}\] right?

OpenStudy (sidsiddhartha):

\[\Large z=[\cos(2n+1)(\pi/4)+isin(2n+1)(\pi/4)]\]

OpenStudy (sidsiddhartha):

so for four poles i have to replace n=0,1,2,3 right?

OpenStudy (anonymous):

Yes that's right. The beauty of the contour you're using (I'm assuming a semicircle in the upper half-plane?) is that you need only consider the residues of the poles in the upper half-plane.

OpenStudy (anonymous):